The hydrogen spectrum includes a red line at 656 nm and a blue-violet line at 434 nm. If light from a hydrogen lamp is incident on a diffraction grating that has 4500 groove/cm, what is the distance between the 2nd order maxima for the red and blue-violet lines on the same side of the central maximum that is imaged on a large screen that is 1.50 m away?
Answer #1

Slit width, d= \frac{1}{4500* 10^2} = 2.22 * 10^{-6} m

Condition for maximum : y_m = \frac{m \lambda D}{d}

For  red line and a blue-violet line,  distance between the 2nd order maxima is

x = \frac{m \lambda_{Red} D}{d} - \frac{m \lambda_{Blue-Violet} D}{d} = \frac{m D}{d} ( \lambda_{Red} - \lambda_{Blue-Violet} )

x = \frac{2*1.5}{2.22 * 10^{-6}} ( 656 - 434 )* 10^{-9} = 0.3 m